OpenStudy (anonymous):

How many coulombs could be in a lightning bolt? A. 1. 6 × 10^19 C B. 5.0 × 10^6 C C. 5.0 × 10^0 C D. 6.24 × 10^18 C

OpenStudy (michele_laino):

we have to know the current carried by a lightning

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

how do we find that out? :/

OpenStudy (michele_laino):

tipically we have: current I = 2*10^5 amperes for t= 1.5 seconds, so the electricity carried by a ligthning is: Q=2*10^5* 1.5 = 4 *10 ^5 Coulombs

OpenStudy (anonymous):

ok! so would our solution be choice B? :/

OpenStudy (michele_laino):

yes! I think so!

OpenStudy (anonymous):

yay! thank you!!:)

OpenStudy (michele_laino):

:)

OpenStudy (anonymous):

oh sorry!! in the answer choices, i forgot to put the negative signs! here are the correct answer choices: A. 1. 6 × 10^-19 C B. 5.0 × 10^-6 C C. 5.0 × 10^0 C D. 6.24 × 10^18 C would this still be choice B? @Michele_Laino :)

OpenStudy (michele_laino):

option D is a very large quantity of electricity

OpenStudy (anonymous):

ohhh okay! which would it be then? :/ i am not sure :P

OpenStudy (michele_laino):

we can conjecture a series of discharges, so I think option D

OpenStudy (michele_laino):

please wait

OpenStudy (anonymous):

ok!

OpenStudy (michele_laino):

if we conjecture a current I = 2*10^3 amperes, and a time interval t= 5*10^(-3) seconds, then Q=2*10^3*5*10^-3=10 coulombs

OpenStudy (anonymous):

so that would mean choice C? :/

OpenStudy (michele_laino):

all data are from Wikipedia

OpenStudy (anonymous):

ah okie :P wikipedia has everything! haha :P

OpenStudy (michele_laino):

yes! I think option C

OpenStudy (anonymous):

okie yay!! again, sorry about this error!! thank you!!

OpenStudy (michele_laino):

option D is too high value

OpenStudy (anonymous):

ahh yes:) okie!! :) back to the other one i posted:P haha sorry again!! :P

OpenStudy (michele_laino):

:)